OpenStudy (anonymous):

What is the probability that you win the California Super Lotto Plus? The California super lotto is played by selecting ve numbers between 1 and 47, and one super ball between the numbers 1 and 27.

OpenStudy (anonymous):

what are ve numbers?

jimthompson5910 (jim_thompson5910):

probably meant to say "five"

OpenStudy (anonymous):

oooh i have seen people use ve to mean negative, never knew why

jimthompson5910 (jim_thompson5910):

oh true, +ve = positive -ve = negative

jimthompson5910 (jim_thompson5910):

I don't like that notation though

OpenStudy (anonymous):

Sorry, yes I meant "five" @satellite73 @jim_thompson5910

OpenStudy (anonymous):

is this a statistic problem? @jim_thompson5910 @satellite73

jimthompson5910 (jim_thompson5910):

close, a probability problem

jimthompson5910 (jim_thompson5910):

how many ways are there to pick the 5 numbers?

OpenStudy (anonymous):

well if there are 47 numbers. so I multiply 47X5? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

more like, 47^5 = 47*47*47*47*47 = ??

OpenStudy (anonymous):

thats 229345007 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

then you multiply that by 27 because there are 27 ways to pick the 6th number

OpenStudy (anonymous):

6192315189

OpenStudy (anonymous):

@jim_thompson5910 ^

jimthompson5910 (jim_thompson5910):

looks good

jimthompson5910 (jim_thompson5910):

the probability of picking all 6 numbers correctly is therefore \[\Large \frac{1}{6,192,315,189}\] because there is only one way to pick all 6 numbers correctly

jimthompson5910 (jim_thompson5910):

you can use a calculator to approximate 1/(6,192,315,189) to get 0.0000000001615 a very very small probability

OpenStudy (anonymous):

Thats a very very small chance